OpenStudy (anonymous):

[Sin^2 22 + Sin^2 68 + sin 63 + cos 63 x sin 27] Cos^2 22 + Cos^2 68

OpenStudy (anonymous):

\[ \left[\begin{matrix}\sin^2 22 + \sin^2 68 \\ \cos^2 22 + \cos^268\end{matrix}\right] \ + sin 63 + cos 63. sin 27 ]

OpenStudy (arindameducationusc):

??